Calculate the current flowing in resistor R1 (in mA) and calculate the power dissipated in the resistor. When voltage = 25V and Resistance=560 Ohms
1
Expert's answer
2019-10-22T10:02:07-0400

"Current, I =\frac{V}{R}=\frac{25}{560}=0.04464\ A=44.64\ mA"


"power\ dissipated=I^2R=0.04464^2\times 560=1.116\ W"